OpenStudy (anonymous):

solve the equation for 0

OpenStudy (anonymous):

anyone????

OpenStudy (anonymous):

tan^2(x) = sec^2(x)-1 so the first equation will be 3(sec^2(x)-1) - sec^2(x)-5 = 0. Simplify and you get x = sec^-1(2). x = pi/3. Used wolfram and got an equation for the second and got x=(1/2)(4*pi*n-pi) and the only x that works there is 3pi/2. Are these 2 different problems?

OpenStudy (anonymous):

yeah, their two different problems

OpenStudy (anonymous):

okay, then its pi/3 for top and 3pi/2 for bottom. I believe...

OpenStudy (anonymous):

plug them in and see what they come up with

OpenStudy (anonymous):

alright, gimmie a sec.

OpenStudy (anonymous):

It works.

OpenStudy (anonymous):

howd you get the answer for the 2nd problem..im kinda lost

OpenStudy (anonymous):

I used wolfram for the second. Im sure theres a trig substitution that you use. http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):